# Gradient Dynamics of Attention: How Cross-Entropy Sculpts Bayesian Manifolds

### v2

@@ 9.1. Bayesian Interpretations of Transformers
Several works argue that transformers implement approximate Bayesian inference, either behaviorally or via probing (e.g. Xie et al., 2022; von Oswald et al., 2023). Our companion paper (Agarwal et al., 2025) demonstrates exact Bayesian behavior and geometric signatures in small wind tunnels, while a scaling paper shows similar patterns in production LLMs. The present work explains how gradient dynamics produce these geometries. It further shows that minimum theoretical Bayesian predictive entropy is close to the empirical entropy. This is studied in greater detail in the companion paper.
@@ 9.3. Optimization and Implicit Bias
The responsibility-weighted value updates derived here are reminiscent of neural EM and slot-attention models, where soft assignments drive prototype updates. The key distinction is that, in transformers, responsibilities are computed via content-addressable attention and prototype updates are driven by backpropagated error signals rather than reconstruction likelihoods. Our focus is not on proposing a new EM-style architecture, but on showing how standard cross-entropy training in attention layers induces EM-like specialization dynamics as a consequence of gradient flow.
@@ 10. Limitations and Future Directions
Our toy simulations are intentionally small. Applying the diagnostics in Section Section 8 to full-scale LLM training runs, tracking advantage matrices and manifold formation over time, is a promising direction.
@@ 11. Conclusion
This paper focuses on first-order mechanisms in a minimal setting; companion work establishes that the same geometric structures persist at scale and support exact Bayesian inference in controlled wind tunnels and large models.
Our key findings are:
@@
    Advantage-based routing. Score gradients $\partial L/\partial s_{ij}=\alpha_{ij}(b_{ij}-\mathbb{E}_{\alpha_{i}}[b])$ implement an advantage rule that reallocates attention toward above-average values.
@@
    Responsibility-weighted value updates. Values evolve via $\Delta v_{j}=-\eta\sum_{i}\alpha_{ij}u_{i}$, becoming prototypes for the queries that attend to them.
@@
    Coupled specialization. Routing and content form a positive feedback loop that produces specialized values and focused attention.
@@
Together with our wind-tunnel and scaling papers, this yields a coherent trilogy: optimization dynamics construct Bayesian geometry, and that geometry enables transformers to behave as Bayesian reasoners in context.

### v3

@@ 8.3. Architectural Choices
9. Toward a General Theory of Content-Based Routing
The gradient dynamics derived in Sections 4 and 5 are specific to softmax attention. However, companion work shows that selective state-space models (Mamba) achieve comparable Bayesian inference through a different routing mechanism—input-dependent gating rather than query-key matching. This raises a natural question: is there a unifying principle that explains why both architectures develop Bayesian geometry? We conjecture that the essential structure is content-based value routing, and that any mechanism satisfying this property will exhibit similar gradient dynamics under cross-entropy training.
9.1. Content-Based Routing: An Abstract Definition
Definition 9.1 (Content-Based Value Routing).
A content-based value routing mechanism is a parameterized function that, given a sequence of representations $\{x_{1},\ldots,x_{T}\}$, computes:
  1. (1)
    Routing weights $w_{ij}(\theta;x_{1:T})\in[0,1]$ that depend on the content of positions $i$ and $j$ (not merely their indices), with $\sum_{j}w_{ij}=1$.
  2. (2)
    Value representations $v_{j}(\phi;x_{j})$ computed from input content.
  3. (3)
    Output $o_{i}=\sum_{j}w_{ij}v_{j}$, a routing-weighted combination of values.
The key requirement is that routing weights depend on what is at each position, not just where it is. This excludes:
  • Fixed positional patterns (e.g., “always attend to position t1t-1”)
  • Learned but content-independent weights (e.g., sinusoidal position encodings alone)
  • Architectures where gating is independent of cross-position relationships (e.g., standard LSTM gates)
9.2. Instantiations
Softmax Attention.
The standard transformer attention mechanism is a content-based routing mechanism with:
(45) $w_{ij}$ $=\frac{\exp(q_{i}^{\top}k_{j}/\sqrt{d})}{\sum_{j^{\prime}}\exp(q_{i}^{\top}k_{j^{\prime}}/\sqrt{d})}$
(46) $v_{j}$ $=W_{V}x_{j}$
(47) $o_{i}$ $=\sum_{j}w_{ij}v_{j}$
Routing weights depend on content through the query-key inner product $q_{i}^{\top}k_{j}$.
Selective State-Space Models (Mamba).
Mamba implements content-based routing through input-dependent state-space parameters:
(48) $\Delta_{t},B_{t},C_{t}$ $=\text{proj}(x_{t})\quad\text{(input-dependent)}$
(49) $h_{t}$ $=\exp(A\cdot\Delta_{t})h_{t-1}+\Delta_{t}\cdot B_{t}\cdot x_{t}$
(50) $o_{t}$ $=C_{t}\cdot h_{t}$
The matrices $\Delta_{t}$ (discretization step), $B_{t}$ (input projection), and $C_{t}$ (output projection) are computed from the input $x_{t}$, making the routing content-dependent. The effective routing weight from position $j$ to position $i>j$ is:
$w_{ij}^{\text{SSM}}=C_{i}\cdot\left(\prod_{k=j+1}^{i}\exp(A\cdot\Delta_{k})\right)\cdot\Delta_{j}\cdot B_{j}$
This depends on the content at all intermediate positions, satisfying the content-based criterion.
LSTM (Counterexample).
Standard LSTMs have gates that depend only on the current input and previous hidden state:
(51) $f_{t}$ $=\sigma(W_{f}[h_{t-1},x_{t}])$
(52) $i_{t}$ $=\sigma(W_{i}[h_{t-1},x_{t}])$
Crucially, the forget gate $f_{t}$ and input gate $i_{t}$ do not depend on the relationship between the current position’s content and other positions’ content. The LSTM cannot selectively route based on content similarity across positions—it applies the same gating logic regardless of what information is available elsewhere in the sequence. This explains why LSTMs can accumulate static sufficient statistics (bijection elimination admits a fixed-dimensional statistic) but fail when the statistic must evolve under dynamics (HMM transport) or be indexed by content (binding).
9.3. Conjectured General Gradient Dynamics
We conjecture that any content-based routing mechanism, when trained with cross-entropy loss, exhibits the following gradient structure:
Conjecture 9.2 (Advantage-Based Routing).
Let $w_{ij}(\theta)$ be content-based routing weights parameterized by $\theta$. Under cross-entropy loss $L$, the gradient with respect to routing parameters satisfies:
$\frac{\partial L}{\partial\theta}\propto\sum_{i,j}w_{ij}\cdot\underbrace{(b_{ij}-\mathbb{E}_{w_{i}}[b])}_{\text{advantage}}\cdot\frac{\partial w_{ij}}{\partial\theta}$
where $b_{ij}$ measures the compatibility between the upstream gradient at $i$ and the value at $j$.
Conjecture 9.3 (Responsibility-Weighted Value Updates).
Value parameters $\phi$ evolve according to:
$\frac{\partial L}{\partial\phi}\propto\sum_{i,j}w_{ij}\cdot u_{i}\cdot\frac{\partial v_{j}}{\partial\phi}$
where $u_{i}$ is the upstream gradient. Values are pulled toward the queries/outputs that route to them, weighted by routing strength.
For softmax attention, we proved these conjectures exactly in Section 4. For selective SSMs, the sequential structure complicates the analysis (gradients must backpropagate through the recurrence), but preliminary analysis suggests the same qualitative structure emerges.
9.4. Implications for Bayesian Geometry
If Conjectures 9.2 and 9.3 hold generally, they explain why both attention and selective SSMs develop Bayesian geometry:
  1. (1)
    Routing specialization: Advantage-based updates cause routing to specialize—weights increase for value positions that are “above average” at reducing the loss for each query.
  2. (2)
    Value clustering: Responsibility-weighted updates pull values toward the queries that use them, creating clusters in value space.
  3. (3)
    Manifold formation: The combination produces low-dimensional value manifolds where position encodes task-relevant information (e.g., posterior entropy).
  4. (4)
    EM-like dynamics: The coupled specialization resembles EM: routing weights act as soft responsibilities (E-step), values act as prototypes (M-step).
This explains the empirical finding from Paper 1: transformers and Mamba both achieve Bayesian inference (0.049 and 0.031 bits MAE on HMM tracking), while LSTMs fail (0.416 bits). The LSTM lacks content-based routing and therefore cannot develop the coupled specialization that produces Bayesian geometry.
9.5. Future Work: Formal Analysis of Selective SSMs
A complete theory requires deriving the gradient dynamics for selective SSMs explicitly. The key challenges are:
  1. (1)
    Sequential dependencies: Unlike attention, SSM gradients must backpropagate through the recurrence $h_{t}=f(h_{t-1},x_{t},\Delta_{t},B_{t})$.
  2. (2)
    Coupled parameters: The input-dependent $\Delta$, $B$, $C$ are computed through shared projections, creating coupled dynamics.
  3. (3)
    Continuous vs. discrete: SSMs operate in continuous state-space; the EM analogy may require a continuous relaxation (e.g., variational inference interpretation).
We leave formal derivation for future work, but note that the empirical success of Mamba on Bayesian tasks strongly suggests that similar dynamics emerge. The abstract framework proposed here provides a roadmap for unifying the theory of content-based routing across architectures.
@@ 10.1. Bayesian Interpretations of Transformers
Several works argue that transformers implement approximate Bayesian inference, either behaviorally or via probing (e.g. Xie et al., 2022; von Oswald et al., 2023). Our companion paper (Agarwal et al., 2025) demonstrates exact Bayesian behavior in small wind tunnels, showing that architectures with content-based value routing (transformers and Mamba) succeed while those without it (LSTMs, MLPs) fail. A scaling paper shows similar geometric signatures in production LLMs. The present work explains how gradient dynamics in attention produce these geometries—deriving analogous dynamics for selective state-space models remains future work.
@@ 10.3. Optimization and Implicit Bias
Attention Training Dynamics.
Recent work has analyzed attention optimization from complementary perspectives. Max-margin analyses show directional convergence of attention parameters toward SVM-like solutions. Two-stage dynamics have been identified where early training exhibits parameter condensation followed by key/query activation. Our advantage-based routing perspective complements these findings: the baseline-subtracted compatibility gradient we derive is the mechanism by which attention reallocates mass, and our two-timescale observation (fast routing, slow values) aligns with reports of stage-wise dynamics. The key difference is our focus on the EM-like interpretation and the connection to Bayesian manifold formation. The responsibility-weighted value updates derived here are reminiscent of neural EM and slot-attention models, where soft assignments drive prototype updates. The key distinction is that, in transformers, responsibilities are computed via content-addressable attention and prototype updates are driven by backpropagated error signals rather than reconstruction likelihoods. Our focus is not on proposing a new EM-style architecture, but on showing how standard cross-entropy training in attention layers induces EM-like specialization dynamics as a consequence of gradient flow.
@@ 11. Limitations and Future Directions
Our toy simulations are intentionally small. Applying the diagnostics in Section 8 to full-scale LLM training runs, tracking advantage matrices and manifold formation over time, is a promising direction.
Attention-specific analysis.
Our gradient derivations are specific to softmax attention. Companion work shows that selective state-space models (Mamba) also implement Bayesian inference via content-based routing, but through different mechanics—input-dependent gating rather than query-key matching. Deriving analogous gradient dynamics for selective SSMs, and identifying whether a similar EM-like structure emerges, is an important direction for unifying the theory of content-based routing across architectures.
@@ 12. Conclusion
Paper I establishes what architectures can implement Bayesian inference, taxonomized by three inference primitives: belief accumulation, belief transport, and random-access binding. This paper explains how gradient descent learns to implement these primitives. Our key findings are:
@@
    Advantage-based routing. Score gradients $\partial L/\partial s_{ij}=\alpha_{ij}(b_{ij}-\mathbb{E}_{\alpha_{i}}[b])$ implement an advantage rule that reallocates attention toward above-average values. This dynamic enables random-access binding: the query-key mechanism learns to retrieve by content.
@@
    Responsibility-weighted value updates. Values evolve via $\Delta v_{j}=-\eta\sum_{i}\alpha_{ij}u_{i}$, becoming prototypes for the queries that attend to them. This enables belief accumulation: values aggregate evidence from positions that route to them.
@@
    Coupled specialization. Routing and content form a positive feedback loop that produces specialized values and focused attention. This enables belief transport: content-based routing can track how beliefs evolve through stochastic dynamics.
@@
The EM interpretation explains why the inference primitives emerge: the E-step (attention) learns what information to route where; the M-step (values) learns what beliefs to store. Architectures without content-based routing (LSTMs, MLPs) cannot implement the E-step and therefore cannot learn the coupled specialization that produces Bayesian geometry. LSTMs can accumulate static sufficient statistics through their recurrent state, but their fixed gating cannot implement the content-dependent operations required for transport or binding. Together with Papers I and III, this yields a coherent trilogy instantiating the Structural Theorem on inference primitives and architectural realizability: Paper I establishes existence and separation (Lemma 1); this paper establishes mechanism (Lemma 2)—showing that gradient dynamics reliably construct the primitives an architecture can support; Paper III establishes scaling and completeness (Lemma 3)—showing these mechanisms persist in production LLMs.